OpenStudy (anonymous):

@iGreen A student answered 48 problems on a test correctly and received a grade 96%. How many problems were on the test?

OpenStudy (igreen):

Okay 48 is 96% on the test. So multiplying 0.96 to something gives us 48. Make an equation. 0.96 * x = 48 Solve for 'x'

OpenStudy (anonymous):

x=50

OpenStudy (anonymous):

:\ Wow..

OpenStudy (anonymous):

2,400?

OpenStudy (anonymous):

@SoulEater142 I'm not trying to be mean, but don't answer my questions. You're not helping me at all.

OpenStudy (anonymous):

I think i just gave you the answer princess. -_-

OpenStudy (anonymous):

Okay, but i didn't ask for the answer, i wanted to try it out myself. That's why i'm here for help, not actual answers.

OpenStudy (anonymous):

:/ my bad... most people want answers not more problems and equations.

OpenStudy (anonymous):

It's okay, but that's not how i'm going to learn. I think getting help FOR the answers, is better. :3

OpenStudy (anonymous):

Fineeee

OpenStudy (anonymous):

^-^ Thank you.

OpenStudy (igreen):

@Ms.Parris is an awesome learner \(\Huge\ddot\smile\)

OpenStudy (anonymous):

@iGreen T-Thank you. <3

OpenStudy (igreen):

\(\sf 0.96 \cdot x = 48\) To solve for 'x', we need to get it by itself..0.96 is in the way. 0.96 is being multiplied to 'x', so to get rid of it, we do the opposite, which is..?

OpenStudy (igreen):

Most people on here are just looking for answers..OS is here to help people learn..not to give out free answers..which is why it's against the Code of Conduct..so please refrain from giving direct answers @SoulEater142

OpenStudy (anonymous):

I don't even know how to explain half the stuff to the people who ask the questions... i just know how to answer them myself

OpenStudy (igreen):

Of course, there are always some people who actually want to learn and be a successful person, like @Ms.Parris just showed herself to be.

OpenStudy (anonymous):

@iGreen Do we divide..? O.o

OpenStudy (igreen):

@SoulEater142 If you know how to find the answer on your own, you should be able to explain it to others..tell them the process and let them do it.

OpenStudy (anonymous):

nope... iv'e always been horrible at explaining my answers because i do most of it mentally.

OpenStudy (igreen):

Yes, we divide 0.96 to both sides to get rid of it. \(\sf 0.96 \cdot x = 48\) \(\sf \dfrac{0.96}{0.96} \cdot x = \dfrac{48}{0.96}\) 0.96 divided by itself is equal to 1..so it just cancels out, leaving us with 'x'. \(\sf \cancel{\bcancel{\dfrac{0.96}{0.96}}} \cdot x = \dfrac{48}{0.96}\) \(\sf x = \dfrac{48}{0.96}\)

OpenStudy (anonymous):

Oh okay. YAY <3

OpenStudy (anonymous):

Thank you, iGreen. Are you a real Math teacher? :o

OpenStudy (igreen):

No, I'm just a 14-year-old kid.

OpenStudy (anonymous):

o_o... Cool. :o I'm 13.

OpenStudy (anonymous):

Well if you ever need a straight up answer for something ask me but otherwise ask @iGreen

OpenStudy (igreen):

@SoulEater142 The problem is, "straight up answers" are not allowed here..you can be banned for it..we guide the asker to the solution.

OpenStudy (anonymous):

Then you should see the public questions... 90% of them just get answers. :P

OpenStudy (igreen):

I know..it's a big issue.

OpenStudy (anonymous):

@iGreen I-i gave you a message..